OpenStudy (anonymous):

A system of equations is shown below: 6x-5y = 5 3x + 5y = 4 The x-coordinate of the solution to this system of equations is _____.

OpenStudy (anonymous):

@selfconfidence777

OpenStudy (anonymous):

x=1.

OpenStudy (selfconfidence777):

well you can multiply the top equation by -1 then you'll have: 5y-6x=-5 5y+3x=4 equate them: 6x-5=4-3x solve for 'x'

OpenStudy (selfconfidence777):

@<3bae<3

OpenStudy (anonymous):

Simply add those two equations termwise and equate... you will get the answer..